OpenStudy (anonymous):

Evaluate the following expression. 2^0

OpenStudy (anonymous):

1.

OpenStudy (shamil98):

Anything to power of 0 is 1.

OpenStudy (anonymous):

anything to the power 0 is 1

OpenStudy (anonymous):

what about Evaluate the following expression. 6 -1

OpenStudy (luigi0210):

*Almost anything to the power of 0 is 1, 0^0 is undefined.

OpenStudy (anonymous):

i mean 6^-1

OpenStudy (anonymous):

6^-1 = 1/6

OpenStudy (anonymous):

6^-1= (1/6)^1=1/6

OpenStudy (anonymous):

what about 6^-3

OpenStudy (anonymous):

same rule. \[(1/6)^{3}=1/216\]

OpenStudy (anonymous):

Can someone help me with this lol ? Solve for x: −4|x + 5| = −16 x = 21/4 , x = −11/4 x = −1, x = 9 x = −1, x = −9 No solution